# Relaxed bound on performance of quantum key repeaters and secure content of generic private and independent bits

### 29 Feb 2024 v3

@@ V Mutual information bound for the secure content of a random private bit
This section focuses on random private bits and provides an upper bound on the distillable key. We perform the random choice of a private state differently than it is done in Christandl et al. (2020). There, the private state has been first transformed via a one-way LOCC protocol to the form of the so-called Bell-private bit of the form $\frac{1}{2}\left(|\psi_{+}\rangle\langle\psi_{+}|\otimes\rho_{+}+|\psi_{-}\rangle\langle\psi_{-}|\otimes\rho_{-}\right)$. Then the two states were defined as $\rho_{\pm}:=U\bar{\rho}_{\pm}U$, where $\bar{\rho}_{\pm}$ is arbitrarily chosen state of the form $\bar{\rho}_{+}:=\frac{Q}{TrQ}$ with $Q$ being a projector onto any $d^{2}/2$ dimensional subspace of $A^{\prime}B^{\prime}$ system, and $\bar{\rho}_{-}:=\frac{Q^{\perp}}{TrQ^{\perp}}$. Our approach is different, as it also applies to the smallest dimension of the shield $d_{s}=2$. We do not transform a private bit into a Bell private bit. We rather observe that any private bit is uniquely represented by a (not necessarily normal) operator $X$ of trace norm equal to 1. We then represent this operator as $U\sigma$ and choose a random state $\sigma$ and random unitary $U$ according to the Haar measure.
Let us start by introducing the notion of random private states. Looking at Eq. (9) we have two objects that can be chosen at random: the unitaries $U_{0},U_{1}$ and the shared state $\sigma_{A^{\prime}B^{\prime}}$. Due to the unitary invariance of the ensemble of random quantum states, we can set one of the unitaries $U_{0}=\mathds{1}$ and the second one to be chosen randomly, denoted $U_{1}=U$. For simplicity, let us also introduce the notation $X=U\sigma_{A^{\prime}B^{\prime}}$. Hence, we can rewrite Eq. (9) as
@@
A short introduction into the properties of random unitaries and quantum states can be found in Appendix B. Before statement of the main result, which is the bound on the distillable key of a random private bit, let us first state the following technical fact
@@
From the form of $\gamma_{ABA^{\prime}B^{\prime}}$ we can conclude that half of its eigenvalues are equal to zero. Therefore it suffices to focus on the eigenvalues of the matrix
@@
First we note that for any |$\ket{\psi}$ we have
$\gamma^{\prime}\begin{pmatrix}\ket{\psi}\\ -X^{\dagger}(\sqrt{XX^{\dagger}})^{-1}\ket{\psi}\end{pmatrix}=0.$ (60)
@@
and the desired results follows from direct calculations. ∎
@@
Let $\gamma_{ABA^{\prime}B^{\prime}}$ be a private state defined as in Eq. (57), with the shielding system of dimension $d_{s}\otimes d_{s}$, let also $X=U\sigma_{A^{\prime}B^{\prime}}$, where $U$ is a Haar unitary be an arbitrary and $\rho$ is a random mixed state sampled from the Hilbert-Schmidt distribution. Then as $d\to\infty$ we get
@@
From Propositions 1 and 2 we have that the entropy of a random mixed state $\rho$ of dimension $d_{s}$ asymptotically behaves as $\log_{2}d_{s}-\frac{1}{2\ln 2}-O\left(\frac{\log_{2}d_{s}}{d_{s}}\right)$ we have
$S(AA^{\prime}BB^{\prime})_{\gamma}=2\log_{2}d_{s}-\frac{1}{2\ln 2}-O\left(\frac{\log_{2}d_{s}}{d_{s}^{2}}\right).$ (65)
All is left is to consider the terms $S(AA^{\prime})_{\gamma}$ and $S(BB^{\prime})_{\gamma}$. First we observe that
@@
Putting all of these facts together we obtain the desired result
@@
We have then immediate corollary
@@
For a random private bit $\gamma_{2}$ of arbitrary large dimension of the shield, in the limit of large $d_{s}\rightarrow\infty$, there is
@@
For the private states, the above bound yields a tighter result, than the one given in Theorem 54. Indeed, since $E_{D}^{\rightarrow}(\rho)\geq 0$, the bound of Eq. (54) is at least $2$. However, the one-way repeater rate of a bipartite state $\rho_{AB}$ can not be larger than the distillable key $K_{D}(\rho_{AB})$ Bäuml et al. (2015). This is a trivial bound: any key-repeater protocol can be viewed as a particular LOCC protocol between $A$ and $B$. As such, it can not increase the initial amount of key in the cut $A:B$. We thus have for a generic private state $\gamma$ (not necessarily irreducible) a trivial bound
@@
$0.360674\displaystyle 1+{1\over 4\ln 2}\approx 1+$ 0.360674 . (70)
Let us note here that the bound is far from being small in contrast with the bound for the private state taken at random due to different randomization procedures proposed in Christandl et al. (2020). This is because our technique does not immediately imply that the private state has Hermitian $X$, nor that its positive and negative parts are random states. We believe, however, that the above bound can be made tighter.
@@ VI Localizable private randomness for a generic local independent state
In this section, we study the rates of private randomness that can be distilled from a generic independent bit. We base on the following idea. While (half of the) mutual information is merely a weak bound on the private key. This function reports the exact amount of private randomness content of quantum state in various scenarios Yang et al. (2019).
Let us recall here, that the local independent bit (in bipartite setting) has form
$\alpha_{AA^{\prime}B^{\prime}}=\sum_{i,j=0}^{1}{1\over 2}|i\rangle\langle j|\otimes U_{i}\sigma_{A^{\prime}B^{\prime}}U_{j}^{\dagger}.$ (71)
[…]
@@
We give below the argument for the analysis of the entropy of system $AA^{\prime}$, used in the proof of the next theorem.
@@
We immediately note that the diagonal blocks are partial traces of independent random states. Hence from Nechita et al. (2018) we have that they almost surely converge to the maximally mixed state. Formally we have (only for the first block)
@@
Let $\alpha_{AA^{\prime}B^{\prime}}$ given by Eq. (71) be the randomly generated state by picking up random state $\sigma_{A^{\prime}B^{\prime}}$ and a unitary transformation $U_{1}$ due to Haar measure. Let also $R_{G}:=1+\frac{1}{2\ln 2}+O(\frac{\log d_{s}}{d_{s}^{2}})$. Then the following bounds are achievable.
@@
    $R_{A}\leq R_{G}-o(1)$, when
    $\log d_{s}\geq\frac{1}{2\ln 2}+O(\frac{\log d_{s}}{d_{s}^{2}})-o(1)$ and
    $R_{A}\leq 1+\log d_{s}$ otherwise,
    $R_{B}\leq 1+\frac{1}{2\ln 2}-O(\frac{\log_{2}d_{s}}{d_{s}})$, when
    $\log_{2}d_{s}\geq 1+\frac{1}{2\ln 2}+O(\frac{\log_{2}d_{s}}{d_{s}})$ and
    $R_{B}\leq\log d_{s}$ otherwise.
@@
    $R_{A}\leq R_{G}$ when
    $\log d_{s}\geq\frac{1}{2\ln 2}+O(\frac{\log_{2}d_{s}}{d_{s}^{2}})-o(1)$, and
    $R_{A}\leq 1+\log d_{s}+o(1)$ otherwise.
    $R_{B}\leq R_{G}$ when
    $\log_{2}d_{s}\geq 1+\frac{1}{2\ln 2}+O(\frac{\log d_{s}}{d_{s}})$
    $R_{B}\leq\log_{2}d_{s}-O(\frac{\log d_{s}}{d_{s}})$, otherwise.
@@
  1. 1.
    By Theorem 2, case $1$, the rate of private randomness reads $R_{A}=\log|AA^{\prime}|-\max\{0,S(AA^{\prime}|B^{\prime})\}$. This equals $\log|AA^{\prime}|-\max\{0,S(AA^{\prime}B^{\prime})-S(B^{\prime})\}$. When $S(B^{\prime})>S(AA^{\prime}B^{\prime})$, we have $R_{A}=\log 2d_{s}$, as the analysis is the same as for the private state (see Eq. (66)). On the other hand, if $S(B^{\prime})\leq S(AA^{\prime}B^{\prime})$ which happens for all high enough dimensions $d_{s}$, there is
    $R_{A}$ $=\log 2d_{s}-[(\log d_{s}^{2}-\frac{1}{2\ln 2}$ (77)
    $-O(\frac{\log d_{s}}{d_{s}^{2}}))-(\log d_{s}-o(1))]$
    $=1+\frac{1}{2\ln 2}+O(\frac{\log d_{s}}{d_{s}})-o(1).$ (78)
@@
    In the above we have used the fact, that $S(AA^{\prime}B^{\prime})$ for an independent state equals $S(AA^{\prime}BB^{\prime})$ for a corresponding private bit (i.e. generated by the same twisting $\tau$ and from the same state $\sigma$).
    We have further $R_{B}\leq\log|B^{\prime}|-\max\{0,S(B^{\prime}|AA^{\prime})\}=\log|B^{\prime}|-\max\{0,S(AA^{\prime}B^{\prime})-S(AA^{\prime})\}$. Let us consider first the case when $S(AA^{\prime}B^{\prime})\geq S(AA^{\prime})$ which holds for large enough $d_{s}$ i.e. whenever $\log_{2}d_{s}\geq 1+\frac{1}{2\ln 2}+O(\frac{\log_{2}d_{s}}{d_{s}^{2}})+O(\frac{\log_{2}d_{s}}{d_{s}})$, which is equivalent to $\log_{2}d_{s}\geq 1+\frac{1}{2\ln 2}+O(\frac{\log_{2}d_{s}}{d_{s}})$, due to the big O(.) notation. (If this condition is not met there is clearly $R_{B}=\log d_{s}-O(\frac{\log d_{s}}{d_{s}})$).
    By asymptotic continuity of the von-Neumann entropy Winter (2016) and Proposition 3 we have $|S(AA^{\prime})-\log_{2}2d_{s}|\leq O(\frac{\log_{2}d_{s}}{d_{s}})+O(d_{s}^{-1})=O(\frac{\log_{2}d_{s}}{d_{s}})$ while $S(AA^{\prime}B^{\prime})=(\log_{2}d_{s}^{2}-\frac{1}{2\ln 2}-O(\frac{\log_{2}d_{s}}{d_{s}^{2}}))$ as argued above. To show achievability, we need to use the lower bound: $S(AA^{\prime})\geq\log_{2}2d_{s}-O(\frac{\log_{2}d_{s}}{d_{s}})$. We then obtain that $R_{B}\leq 1+\frac{1}{2\ln 2}-O(\frac{\log_{2}d_{s}}{d_{s}}))\leq\log_{2}|B^{\prime}|-S(B^{\prime}|AA^{\prime})$ is an achievable rate. Here we have included term $O(\frac{\log d_{s}}{d_{s}^{2}}))$ in term $O(\frac{\log_{2}d_{s}}{d_{s}}))$ due to the big $O(.)$ notation.
  2. 2.
    The second case reads the same bounds as the first just without additional conditions on dimension $d_{s}$. This is because in the second case of Theorem 2, the bounds have $S(A|B)$ and $S(B|A)$ terms rather than the $S(A|B)_{+}$ and $S(B|A)_{+}$ in its formulation.
  3. 3.
    The third case stems from an observation, that the global purity reads $R_{G}=\log|AA^{\prime}B^{\prime}|-S(AA^{\prime}B^{\prime})=\log(2d_{s}^{2})-(2\log d_{s}-\frac{1}{2\ln 2}-O(\frac{\log d_{s}}{d_{s}^{2}}))$. The rates $R_{A}$ and $R_{B}$ achive the same bound $R_{G}$ in this case.
  4. 4.
    The last case is due to case $4$ of the Theorem 2: $R_{A}\leq\log|AA^{\prime}B^{\prime}|-\max\{S(B^{\prime}),S(AA^{\prime}B^{\prime})\}$. When $S(B^{\prime})>S(AA^{\prime}B^{\prime})$ i.e. $\log d_{s}<\frac{1}{2\ln 2}+O(\frac{\log_{2}d_{s}}{d_{s}^{2}})-o(1)$, there is $R_{A}\leq 1+2\log_{2}d_{s}-\log_{2}d_{s}+o(1)=1+\log_{2}d_{s}+o(1)$. When $S(B^{\prime})\leq S(AA^{\prime}B^{\prime})$ there is $R_{A}\leq 1+2\log_{2}d_{s}-[2\log_{2}d_{s}-(\frac{1}{2\ln 2}+O(\frac{\log_{2}d_{s}}{d_{s}^{2}})])=1+\frac{1}{2\ln 2}+O(\frac{\log_{2}d_{s}}{d_{s}^{2}})$.
    Finally we have $R_{B}\leq\log|AA^{\prime}B^{\prime}|-\max\{S(AA^{\prime}),S(AA^{\prime}B^{\prime})\}$. When $\log_{2}d_{s}\geq 1+\frac{1}{2\ln 2}+O(\frac{\log d_{s}}{d_{s}})$ we have $S(AA^{\prime})\leq S(AA^{\prime}B^{\prime})$, and then $R_{B}\leq R_{G}=1+2\log d_{s}-(2\log d_{s}-\frac{1}{2\ln 2}-O(\frac{\log d_{s}}{d_{s}^{2}}))=1+\frac{1}{2\ln 2}+O(\frac{\log d_{s}}{d_{s}^{2}})$. Otherwise (if. $S(AA^{\prime})>S(AA^{\prime}B^{\prime})$) there is $R_{B}\leq 1+2\log d_{s}-(\log 2d_{s}+O(\frac{\log d_{s}}{d_{s}}))=\log d_{s}-O(\frac{\log d_{s}}{d_{s}})$.
@@
From the above Theorem we can see that in case of the system $A$, for asymptotically large dimension $d_{s}$ the value of global purity $1+\frac{1}{2\ln 2}+O(\frac{\log d_{s}}{d_{s}^{2}})$ can be reached in all four cases. In the case of system $B$ we invoked only the (lower and upper) bounds on the entropy of involved systems ($AA^{\prime}$), hence the bounds are not tight, possibly less than the maximal achievable ones.

### 09 Mar 2026 v4

@@ V Mutual information bound for the secure content of a random private bit
This section focuses on random private bits and provides an upper bound on the distillable key. We perform the random choice of a private state differently from that in Christandl et al. (2020). There, the private state has been first transformed via a one-way LOCC protocol to the form of the so-called Bell-private bit of the form $\frac{1}{2}\left(|\psi_{+}\rangle\langle\psi_{+}|\otimes\rho_{+}+|\psi_{-}\rangle\langle\psi_{-}|\otimes\rho_{-}\right)$. Then the two states were defined as $\rho_{\pm}:=U\bar{\rho}_{\pm}U$, where $\bar{\rho}_{\pm}$ is arbitrarily chosen state of the form $\bar{\rho}_{+}:=\frac{Q}{TrQ}$ with $Q$ being a projector onto any $d^{2}/2$ dimensional subspace of $A^{\prime}B^{\prime}$ system, and $\bar{\rho}_{-}:=\frac{Q^{\perp}}{TrQ^{\perp}}$ with a projector $Q^{\perp}$ complementary to $Q$. Our approach is different, as it also applies to the smallest dimension of the shield $d_{s}=2$. We do not transform a private bit into a Bell private bit. We rather observe that any private bit is uniquely represented by a (not necessarily normal) operator $X$ of trace norm equal to 1. We then represent this operator as $U\sigma$ and choose a random state $\sigma$ and random unitary $U$ according to the Haar measure.
We begin by introducing the notion of random private states. Looking at Eq. (10), we have two objects that can be chosen at random: the unitaries $U_{0},U_{1}$ and the shared state $\sigma_{A^{\prime}B^{\prime}}$. Due to the unitary invariance of the ensemble of random quantum states, we can set one of the unitaries $U_{0}=\mathds{1}$ and the second one to be chosen randomly, denoted $U_{1}=U$. For simplicity, let us also introduce the notation $X=U\sigma_{A^{\prime}B^{\prime}}$. Hence, we can rewrite Eq. (10) as
@@
A short introduction to the properties of random unitaries and quantum states can be found in Appendix A. Before the statement of the main result, which is the bound on the distillable key of a random private bit, let us first state the following technical fact
@@
From the form of $\gamma_{ABA^{\prime}B^{\prime}}$ we can conclude that half of its eigenvalues are equal to zero. Therefore, it suffices to focus on the eigenvalues of the matrix
@@
[…]
@@
and the desired results follow from direct calculations. ∎
@@
Let $\gamma_{ABA^{\prime}B^{\prime}}$ be a private state defined as in Eq. (78), with the shielding system of dimension $d_{s}\otimes d_{s}$, let also $X=U\sigma_{A^{\prime}B^{\prime}}$, where $U$ is a Haar unitary be an arbitrary and $\sigma_{A^{\prime}B^{\prime}}$ is a random mixed state sampled from the Hilbert-Schmidt distribution. Then as $d\to\infty$ we get
@@
We first justify the entropy of the full state $\gamma$. Recall that $X=U\sigma_{A^{\prime}B^{\prime}}$ with $\sigma_{A^{\prime}B^{\prime}}\geq 0$ and $\operatorname{Tr}\sigma_{A^{\prime}B^{\prime}}=1$. Then
$\sqrt{XX^{\dagger}}=\sqrt{U\sigma_{A^{\prime}B^{\prime}}^{2}U^{\dagger}}=U\sigma_{A^{\prime}B^{\prime}}U^{\dagger},$ (86)
hence $\sqrt{XX^{\dagger}}$ and $\sigma_{A^{\prime}B^{\prime}}$ have identical spectra. By Proposition 1, the non-zero eigenvalues of $\gamma$ coincide with those of $\sqrt{XX^{\dagger}}$ (the remaining eigenvalues are zeros, which do not contribute to the von-Neumann entropy). Therefore
$S(AA^{\prime}BB^{\prime})_{\gamma}=S(\gamma)=S(\sqrt{XX^{\dagger}})=S(\sigma_{A^{\prime}B^{\prime}}).$ (87)
Applying Proposition 2 with $d=d_{s}^{2}$ yields
$S(AA^{\prime}BB^{\prime})_{\gamma}=\log_{2}d_{s}^{2}-\frac{1}{2\ln 2}-O\left(\frac{\log_{2}d_{s}^{2}}{d_{s}^{2}}\right).$ (88)
All is left is to consider the terms $S(AA^{\prime})_{\gamma}$ and $S(BB^{\prime})_{\gamma}$. First, we observe that
@@
Putting all of these facts together, we obtain the desired result
@@
We then have the immediate corollary.
@@
For a random private bit $\gamma_{2}$ of an arbitrarily large dimension of the shield, in the limit of large $d_{s}\rightarrow\infty$, there is
@@
For the private states, the above bound yields an independent result than the one given in Theorem 4. Indeed, the one-way repeater rate of a bipartite state $\rho_{AB}$ can not be larger than the distillable key $K_{D}(\rho_{AB})$ Bäuml et al. (2015). This is a trivial bound: any key-repeater protocol can be viewed as a particular LOCC protocol between $A$ and $B$. As such, it can not increase the initial amount of key in the cut $A:B$. We thus have for a generic private state $\gamma$ (not necessarily irreducible) a trivial bound
@@
$1+{1\over 4\ln 2}\approx 1+0.360674.$ (93)
Let us note here that the bound is far from being small, in contrast to the bound for the private state taken at random, due to the different randomization procedures proposed in Christandl et al. (2020). This is because our technique does not immediately imply that the private state has Hermitian $X$, nor that its positive and negative parts are random states. We believe, however, that the above bound can be made tighter.
@@ VI Localizable private randomness for a generic local independent state
In this section, we study the rates of private randomness that can be distilled from a generic independent bit. We base our results on the following idea. While (half of the) mutual information is merely a weak bound on the private key, this function reports the exact amount of private randomness content of a quantum state in various scenarios Yang et al. (2019).
Let us recall here that the local independent bit (in a bipartite setting) has the form
$\alpha_{AA^{\prime}B^{\prime}}=\sum_{i,j=0}^{1}{1\over 2}|i\rangle\langle j|_{A}\otimes U_{i}\sigma_{A^{\prime}B^{\prime}}U_{j}^{\dagger}.$ (94)
where $U_{i}$ are unitary transformations acting on system $A^{\prime}B^{\prime}$ and $\sigma$ is an arbitrary state on the latter system. We assume that the system $A^{\prime}B^{\prime}$ is of dimension $d_{s}\otimes d_{s}$. In the above we borrow the notation of Yang et al. (2019) where $\alpha$ denotes the independent states, thus $\alpha$ should not be confused with the parameter of the sandwiched Rényi relative entropy given in Eq. (25). In the matrix form, it is as follows
@@
We immediately note that the diagonal blocks are partial traces of independent random states. Hence, by Nechita et al. (2018), they almost surely converge to the maximally mixed state. Formally, we have (only for the first block)
@@
$\mathbb{E}_{U_{1}}\|B\|_{2}^{2}=\frac{1}{d}\operatorname{Tr}(\sigma_{A^{\prime}B^{\prime}}^{2}).$ (100)
For $\sigma_{A^{\prime}B^{\prime}}$ sampled from the Hilbert–Schmidt distribution on $\mathbb{C}^{d^{2}}$ we have $\mathbb{E}\operatorname{Tr}(\sigma_{A^{\prime}B^{\prime}}^{2})=O(d^{-2})$ (see, e.g., induced-state moment formulas in Życzkowski and Sommers (2001)), hence $\mathbb{E}\|B\|_{2}^{2}=O(d^{-3})$ and therefore $\|B\|_{\infty}\leq\|B\|_{2}=O(d^{-3/2})$. Consequently $d\|B\|_{\infty}=O(d^{-1/2})\to 0$.
The diagonal blocks are reduced states of independent Hilbert–Schmidt random states; by concentration for induced states (see Nechita et al. (2018)),
[…]
almost surely. Putting diagonal and off-diagonal bounds together yields
@@
Let $\alpha_{AA^{\prime}B^{\prime}}$ given by Eq. (94) be the randomly generated state by picking up a random state $\sigma_{A^{\prime}B^{\prime}}$ and a unitary transformation $U_{1}$ due to Haar measure. Let also $R_{G}:=1+\frac{1}{2\ln 2}$. Then for all but finitely many $d_{s}\in\mathbb{N}$ the following bounds are achievable
@@
    $R_{A}<R_{G}$, $R_{B}<\frac{1}{2\ln 2}$,
    $R_{A}+R_{B}\leq R_{G}$,
@@
    $R_{A},R_{B}\leq R_{G}$, $R_{A}+R_{B}\leq R_{G}$.
@@
At first observe that for the considered state $\alpha_{AA^{\prime}B^{\prime}}$, we have
$\log_{2}|AA^{\prime}B^{\prime}|-S(AA^{\prime}B^{\prime})=$ (103)
$\log_{2}2d_{s}^{2}-\log_{2}d_{s}^{2}+\frac{1}{2\ln 2}+O\left(\frac{\log_{2}d_{s}^{2}}{d_{s}^{2}}\right).$ (104)
Thus, its global purity satisfies
$\log_{2}|AA^{\prime}B^{\prime}|-S(AA^{\prime}B^{\prime})\geq 1+\frac{1}{2\ln 2}.$ (105)
@@
In the above we have used Equation (88) and the fact, that $S(AA^{\prime}B^{\prime})$ for an independent state $\alpha_{AA^{\prime}B^{\prime}}$ equals $S(AA^{\prime}BB^{\prime})$ for a corresponding private bit (i.e., generated by the same twisting $\tau$ and from the same state $\sigma$). Hence, $R_{G}=1+\frac{1}{2\ln 2}$ lowerbounds global purity of the considered $\alpha_{AA^{\prime}B^{\prime}}$ state. We are now ready to prove each case of the theorem separately.
1. By Theorem 2, case $1$, the necessary condition for the rate of private randomness $R_{A}$ to be achievable, is $R_{A}\leq\log_{2}|AA^{\prime}|-\max\{0,S(AA^{\prime}|B^{\prime})\}$. We want to simplify this formula. Observe that
$S(AA^{\prime}|B^{\prime})=S(AA^{\prime}B^{\prime})-S(B^{\prime})\geq$ (106)
$\log_{2}d_{s}-\frac{1}{2\ln 2}-C_{AA^{\prime}B^{\prime}}\frac{\log_{2}d_{s}^{2}}{d_{s}^{2}},$ (107)
for some constant $C_{AA^{\prime}B^{\prime}}$. In the above, we have used equations (88) and (90) to bound terms $S(AA^{\prime}B^{\prime})$ and $S(B^{\prime})$ respectively. Surely, $\log_{2}d_{s}-\frac{1}{2\ln 2}-C_{AA^{\prime}B^{\prime}}\frac{\log_{2}d_{s}^{2}}{d_{s}^{2}}\geq 0$ for all but finitely many $d_{s}$. Thus we can consider only the case in which $\max\{0,S(AA^{\prime}|B^{\prime})\}=S(AA^{\prime}|B^{\prime})$. Now, we aim to upperbound the $S(AA^{\prime}|B^{\prime})$ term. Notice that
$S(AA^{\prime}|B^{\prime})=S(AA^{\prime}B^{\prime})-S(B^{\prime})\leq$ (108)
$\log_{2}d_{s}^{2}-\frac{1}{2\ln 2}-\log_{2}d_{s}+f_{B^{\prime}}(d_{s})=$ (109)
$\log_{2}d_{s}-\frac{1}{2\ln 2}+f_{B^{\prime}}(d_{s}),$ (110)
for $f_{B^{\prime}}(d_{s})=o(1)$. Hence, we have
$\log_{2}|AA^{\prime}|-\max\{0,S(AA^{\prime}|B^{\prime})\}\geq$ (111)
$\log_{2}2d_{s}-\log_{2}d_{s}+\frac{1}{2\ln 2}-f_{B^{\prime}}(d_{s})=$ (112)
$1+\frac{1}{2\ln 2}-f_{B^{\prime}}(d_{s}),$ (113)
so the rate $R_{A}$ satisfying $R_{A}\leq 1+\frac{1}{2\ln 2}-f_{B^{\prime}}(d_{s})$ is achievable. This implies that all rates satisfying $R_{A}<1+\frac{1}{2\ln 2}$ are achievable for all but finitely many $d_{s}$. On the other hand, by analogous arguments, one can show that all rates $R_{B}$ satisfying
$R_{B}\leq\frac{1}{2\ln 2}-f_{AA^{\prime}}(d_{s}),$ (114)
where $f_{AA^{\prime}}(d_{s})=o(1)$, are achievable. Once again, this implies that all rates $R_{B}$ satisfying $R_{B}<1+\frac{1}{2\ln 2}$ are achievable for all but finitely many $d_{s}$. Finally, notice that the pair $(R_{A},R_{B})$ is achievable only if the constraint $R_{A}+R_{B}\leq R_{G}$ is satisfied.
2. Proof of this case is the same as for case 1.
3. Follows directly from Theorem 2 case 3.
4. For all but finitely many $d_{s}$, inequalities $S(AA^{\prime}B^{\prime})\geq S(AA^{\prime})$ and $S(AA^{\prime}B^{\prime})\geq S(B^{\prime})$ are satisfied. Hence, formulas from Theorem 2 case 4 for achievable rates $R_{A}$ and $R_{B}$, reduce to $\log_{2}|AA^{\prime}B^{\prime}|-S(AA^{\prime}B^{\prime})$. As we showed previously, $\log_{2}|AA^{\prime}B^{\prime}|-S(AA^{\prime}B^{\prime})\geq 1+\frac{1}{2\ln 2}=R_{G}$. Thus, for all but finitely many $d_{s}$ the rates satisfying $R_{A},R_{B}\leq R_{G}$, $R_{A}+R_{B}\leq R_{G}$ are achievable. ∎
@@
From the above theorem, we can see that in case of the system $A$, for asymptotically large dimension $d_{s}$ the value $1+\frac{1}{2\ln 2}$ can be approached in all four cases. In the case of system $B$ we invoked only the (lower and upper) bounds on the entropy of involved systems ($AA^{\prime}$). Hence, the bounds are not tight, possibly less than the maximal achievable ones.